2. Bombing Raid. During World War lI, Allied forces carried out daring, daylight air raids to bomb targets in Nazi Germany. (This was the inspiration for the 1990 film Memphis Belle.) Imagine you are the bombardier (person in charge of dropping the bombs) in a B-17 bomber flying toward your target at an altitude of 7000 m at a speed of 300 km/h. (As always, in our imaginary world there is no air resistance.) a. How many minutes before passing over the target should you drop your bombs? b. As bombardier, you look down through a bombsight to line the plane up with your target. At what angle with respect to vertical should your bombsight be looking to ensure that your target is in the sights at the time you must drop your bombs?
Answer #1

a]time taken by target to reach ground t is given by second equation of motion, h = 0.5gt^2

t = sqrt(2h/g) = sqrt(2*7000/9.8) = 37.8 s = 0.63 min

minutes before = 0.63 minutes answer

b] horizontal distance d =vt= 300*0.63/60 = 3.15 km = 3150 m

angle = arctan(3150/7000) = 24.23 degree answer